OpenStudy (anonymous):

for which integer values of "x" x^2+19x+92 is a perfect square. I want to know the procedure

hartnn (hartnn):

\(ax^2+bx+c\) will become a perfect square if it has same (equal) roots, like example : if the roots were 2,2 , then it'd be a perfect square..

hartnn (hartnn):

and for \(ax^2+bx+c \) to be a perfect square, its discriminant must be 0 you know what a discriminant of a quadratic expression is ?

OpenStudy (anonymous):

yes I know about d discriminant but how can we say that roots have to be equal for it to be a perfect square?

hartnn (hartnn):

i must have mis-read the question....let me think

OpenStudy (anonymous):

ok.

hartnn (hartnn):

do you know how to complete the square ?

OpenStudy (anonymous):

yes.

OpenStudy (2bornot2b):

If the roots are equal then you end up with something like (x-a)(x-a) after factorizing your equation, where 'a' is the root.

OpenStudy (2bornot2b):

So that is why you must have repeated roots

hartnn (hartnn):

can you complete the square for that quadratic expression for me ?

OpenStudy (anonymous):

yes.give me a min.

OpenStudy (anonymous):

[x+(19/2)]^2 + 3/2

OpenStudy (anonymous):

start with x^2+19x+92=n^2

OpenStudy (anonymous):

I tried that but I am not getting what to do with it. @BSwan

OpenStudy (anonymous):

its number theory :)

OpenStudy (anonymous):

why gave medal :-\ i still dint solve it yet

OpenStudy (anonymous):

do you what formula the perfect square should have ?

OpenStudy (anonymous):

I gave u the medal coz I got the solution thru ur approach.

OpenStudy (anonymous):

could you plz share ur sol ? cuz i found one too

OpenStudy (anonymous):

my hint is :- the square of any integer is either of the form 3k, or 3k+11

OpenStudy (anonymous):

ok lol fair enough :)

OpenStudy (anonymous):

tell me the reason behind ur hint,i.e.its logic @BSwan

OpenStudy (anonymous):

its using the division algorithm

OpenStudy (anonymous):

plz elaborate

OpenStudy (anonymous):

so let x=3n,3n+1,3n+2 x^2+19x+92 case 1 :- (3n)^2+19(3n)+92 9n^2+57n+92=3(3n^2+19n)+2 (mod 3) of the form 3k+2 something like this , but i cant remember xD so ignore it

OpenStudy (anonymous):

ok.